OpenStudy (anonymous):

Find all solutions in the interval 0

OpenStudy (anonymous):

I first divide the right by 6 making it 1/2 then take the square root of both sides getting me cos(theta)= sqrt(2)/2....

OpenStudy (anonymous):

but then thats when i get confused

OpenStudy (***[isuru]***):

hi, \[6 \cos ^{2} \theta = 3\] \[\cos ^{2}\theta = \frac{ 3 }{6 } = \frac{ 1 }{2 }\] \[\cos \theta = \pm \sqrt{\frac{ 1 }{ 2 }}\] \[\cos \theta = \pm \frac{ 1 }{\sqrt{2} }\] when \[\cos \theta = \frac{ 1 }{ \sqrt{2} }\] \[\cos \theta = \cos \frac{ \pi }{ 4 }\] \[\theta = 2n \pi \pm \frac{ \pi }{4 } where \ \left( n \epsilon \mathbb{Z} \right)\] when \[\cos \theta = - \frac{ 1 }{ \sqrt{2} }\] \[\cos \theta = \cos( \pi - \frac{ \pi }{ 4 })\] \[\theta = 2n \pi \pm ( \pi - \frac{ \pi }{ 4 }) where \ \left( n \epsilon \mathbb{Z} \right)\]

OpenStudy (***[isuru]***):

did ya get it... ? srry it takes time to type the equations :/

OpenStudy (anonymous):

thank you for your detailed explanation.. but isnt 1/sqrt2 just sqrt2/2?...

OpenStudy (anonymous):

I've always understood that you can never have a sqrt at the bottom. maybe im wrong

OpenStudy (***[isuru]***):

yeah!! they r same....

OpenStudy (anonymous):

ahhh ok, big big help thanks man!!!! i really appreciate it.

OpenStudy (anonymous):

Rvenegas13=Fan lmfao

OpenStudy (***[isuru]***):

but don't u know that \[\cos \frac{ \pi }{ 4} = \frac{ 1 }{ \sqrt{2} } = \frac{ \sqrt{2} }{2 }\]

OpenStudy (anonymous):

which is 45deg

OpenStudy (***[isuru]***):

yep!!